Suppose that 10 people live on a street and that each of them is willing to pay GHS 2 for each extra streetlight, regardless of the number of streetlights provided. If the cost of providing x streetlights is given by 2 )( x xC  , what is the Pareto efficient number of streetlights to provide?
1
Expert's answer
2020-07-03T11:05:02-0400

If the cost of providing x streetlights is given by C(x) = x^2, then the Pareto efficient number of streetlights to provide is:

"10\times2\times x = x^2,"

x = 20 streetlights.